# Comparing the latent features of universal machine-learning interatomic potentials

### v1

@@ VII.3.2 PET-OMAT
From the error heat maps (Fig. S6), the models share a common backbone in the latent space: smaller models (XS, S) can capture the main latent feature manifold of larger models (M, L) with reasonable accuracy. Local reconstruction errors are higher and have stronger asymmetry, which indicates that local feature neighborhoods are progressively better reconstructed as the model size increases. Notably, the M model provides a favorable tradeoff, achieving low errors when reconstructing the L model, and can thus serve as an effective pre-training base for fine-tuning while remaining computationally efficient.
Figure S6: Reconstruction errors among PET-OMAT checkpoints of different sizes on the MAD test dataset.
VII.4 Single vs. Multi-head architecture of the same model
We next compare latent spaces between the single-head MACE-OMAT-0, original MACE architecture MACE fully trained on OMat24 omat24 (MACE-OMAT-0) against the corresponding OMat-specific readout head of the multi-head MACE-MH-1 model mace_multihead. The multi-head variant is the original architecture enhanced with nonlinear interaction blocks and trained with multi-head replay fine-tuning, where task-specific readout heads are attached to a shared backbone and simultaneously trained on new data plus a replay buffer sampled from the original foundational training data, to prevent catastrophic forgetting of the pre-trained representations.
As shown in Table S4, reconstruction is nearly symmetric with moderate GFRE but very low LFRE. This suggests that the multi-head design preserves core representational fidelity while improving the global readout, as intended by the multi-head replay fine-tuning. The latent feature space remains essentially unchanged, while the readout becomes more expressive. These results also could validate that models trained with the new multi-head protocol are highly backward-compatible in representation space with earlier single-head MACE checkpoints.
@@ VIII Varying targets
We evaluated the information content overlap in the last-layer features of models between PET-MAD (trained on energies and forces) and PET-MAD-DOS (trained on the electronic density of states) by computing reconstruction errors on the MAD test dataset. The results show lower errors when reconstructing PET-MAD features from PET-MAD-DOS features compared to the reverse direction (Table S5). It indicates that the DOS-based features capture aspects sufficient for approximating energy-based features, while including additional information not fully aligned with the energy-based latent feature space.
IX Fine-tuning
IX.1 Training errors
Each PET checkpoint corresponding to different training strategies was evaluated on the test split of the LPS dataset, comprising 412 structures, on MAE and RMSE on energies and forces. Results are summarized in Table S6.
@@ IX.1 Training errors
IX.2 Hyperparameters
We report the hyperparameters used for all PET-based models trained on LPS across the training strategies described in Sec. III.4.
@@ IX.2 Hyperparameters
IX.3 Latent space projection
We apply PCovR projection on the last-layer features extracted from PET checkpoints obtained with different fine-tuning techniques: full fine-tuning (FT), head fine-tuning (HF), full transfer learning (FTL), and head transfer learning (HTL) upon the LPS dataset. We also extract features from the basic PET-MAD uMLIP, used as a pre-training basis. For the regression term of PCovR, we use the ground-truth cohesive energy, and we form the structure-level features by concatenating the mean and standard deviation of the atomic feature vectors. We set the mixing parameter $\alpha=0.5$ to balance the regression and PCA terms, thereby enforcing consistency between checkpoint projections. The PCovR latent space was first fitted using the embeddings on the MAD test subset, and the LPS projections were then projected into this MAD latent space.
The projections are consistent and similar across checkpoints, meaning that PET-derived representations retain the same global structure due to the common pre-trained backbone. The fluctuations are mostly related to the transfer learning variants (FTL, HTL), which means stronger local adaptations that change how LPS structures are associated with regions of the MAD space.
Figure S7: Two-dimensional PCovR projections of the LPS dataset using last-layer features of PET checkpoints trained with different training setups. The greyscale background points correspond to the test subset of the MAD dataset, and the colored points in each panel to the PCovR projections of the LPS into the MAD latent space.
The similarity and small fluctuations of the projections align with the minimal global and local reconstruction errors described in Section III.1.
X Backbone vs. last-layer features
We compare backbone (BB) and last-layer (LL) features across five PET variants, described in SI VII.1.2. Checkpoints pre-trained on OMat24, corresponding to OAM, OMAD, OMAT, and OMATPES columns, show higher LL to BB reconstruction errors (Fig. S8) than the non-pre-trained MAD, which might indicate that pre-training enriches BB with transferable information. Overall, the flow is asymmetric, with GFRE averages 0.29 for LL $\rightarrow$ BB and 0.19 for BB $\rightarrow$ LL, suggesting that BB features are richer with the MLP head discarding details, e.g., non-invariant components, to suit target prediction.
Figure S8: Global and local reconstruction errors for last-layer features ($L$) and backbone features ($B$) among different PET checkpoints taken on the MAD test dataset.
XI Local to global features
@@
Figure S9: Reconstruction errors of the eight-order cumulative structure-level features among uMLIPs on the MAD test dataset.

### v2

@@ VIII.3.2 PET-OMAT
From the error heat maps (see Fig. S7), the models seem to share a common backbone in the latent space: smaller models (XS, S) can capture the main latent feature manifold of larger models (M, L) with reasonable accuracy. Local reconstruction errors are higher and have stronger asymmetry, which indicates that local feature neighborhoods are progressively better reconstructed as the model size increases. We argue that the M model provides a favorable tradeoff, achieving low errors when reconstructing the L model, and can thus serve as an effective pre-training base for fine-tuning while remaining computationally efficient.
Figure S7: Reconstruction errors across the PET-OMAT checkpoints of different sizes computed for the MAD test subset.
VIII.4 Single vs. multi-head architecture of the same model
We next compare the latent spaces of the single-head MACE-OMAT-0, the MACE architecture Batatia et al. (2025a) fully trained on OMat24 Barroso-Luque et al. (2024) (MACE-OMAT-0), against the corresponding OMat24-specific readout head of the multi-head MACE-MH-1 model Batatia et al. (2025b). The multi-head variant is the MACE architecture enhanced with nonlinear interaction blocks and trained with multi-head replay fine-tuning, where task-specific readout heads are attached to a shared backbone and simultaneously trained on new data plus a replay buffer sampled from the original foundational training data to prevent catastrophic forgetting. As shown in Table S4, reconstruction is nearly symmetric with moderate GFRE but very low LFRE. This hints that the multi-head design preserves the local representational fidelity while enhancing the global readout. In other words, the latent feature space remains essentially unchanged locally, but globally, the newly devised training strategy allows the multi-head model to alter the manifold.
@@ IX Varying targets
Here, we evaluate the information content overlap between the last-layer features of PET-MAD (trained on energies and forces) and PET-MAD-DOS (trained on the electronic density of states) via feature reconstruction errors on the MAD test subset. The results show lower errors when reconstructing PET-MAD features from PET-MAD-DOS features compared to the reverse direction (Table S5). It indicates that the DOS-based features capture aspects sufficient for approximating energy-based features, while encoding additional information that the energy-only latent feature space may be oblivious to.
X Fine-tuning
X.1 Training errors and hyperparameters
The errors of each PET checkpoints corresponding to the different fine-tuning strategies have been evaluated on the test split of the LPS dataset composed of 412 structures. Results are shown in Table S6. We also report the hyperparameters used for the PET-based models in Table S7.
@@ X.1 Training errors and hyperparameters
X.2 Latent space projection
In addition to the analyses in the main text, here we perform the PCovR analysis using the last-layer features extracted from PET checkpoints obtained with different fine-tuning techniques: full fine-tuning (FT), head fine-tuning (HF), full transfer learning (FTL), and head transfer learning (HTL) upon the LPS dataset. We also extract corresponding features from PET-MAD, which has been used as a pre-training basis. For the regression term of PCovR, we use the ground-truth cohesive energy, and we form the structure-level features by using the second order cumulant or the concatenation of the mean and standard deviation of the atomic feature vectors. We set the mixing parameter $\alpha=0.5$ to balance the regression and PCA terms, thereby enforcing consistency between checkpoint projections. Note that the PCovR space was first fitted using the embeddings from the MAD test subset, and the LPS projections were simply projected into this space.
The projections are consistent and similar across checkpoints, meaning that PET-derived representations retain the same global structure due to the common pre-trained backbone. We note that relatively larger fluctuations are observed for the transfer learning variants (FTL, HTL), which can be explained by the involvement of new heads with randomized initial weights. The similarities and differences observed in the PCovR projects are well-aligned with the global and local feature reconstruction errors presented in Section III.1.
Figure S8: Two-dimensional PCovR projections of the LPS dataset constructed using the last-layer features of PET checkpoints trained with different fine-tuning setups. The greyscale background points correspond to the test subset of the MAD dataset, and the colored points in each panel correspond to the PCovR projections of the LPS systems “within” the MAD latent space.
XI Backbone vs. last-layer features
We compare backbone (BB) and last-layer (LL) features across five PET variants, described in SI VIII.1.2. Checkpoints pre-trained on OMat24, corresponding to OAM, OMAD, OMAT, and OMATPES columns, show higher LL to BB reconstruction errors (Fig. S9) than the non-pre-trained MAD, which indicates that pre-training can enrich the BB features with transferable information. Overall, the flow is asymmetric, with an average GFREs of 0.29 for LL $\rightarrow$ BB and 0.19 for BB $\rightarrow$ LL, suggesting that BB features encode more information compared to the LL features.
Figure S9: Global and local reconstruction errors for the last-layer features ($L$) and backbone features ($B$) of different PET checkpoints computed for the MAD test dataset.
XII Local to global features
@@
The global and local feature reconstruction errors across the uMLIPs computed with $\boldsymbol{\kappa}^{(8)}_{S}$ are presented in Fig. S10.
Figure S10: Reconstruction errors of the eight-order cumulative structure-level features across the uMLIPs on the MAD test dataset.